Topic: Word problem two-fifth of a number subtracted from one-quarter of itself is equal to -3.What is the number?
Answer by stanbon(75887)

Word problem: two-fifth of a number subtracted from one-quarter of itself is equal to -3. What is the number?
----
(1/4)x - (2/5)x = -3
------
[(1/4)-(2/5)]x = -3
-----
[(5-8)/20]x = -3
-----
(-3/20)x = -3
(1/20)x = 1
x = 20
-------------
Cheers,
Stan H.
==============